OpenStudy (anonymous):

Q5) For the data Vs=5V, R=7.5 kohm k=3(mA/√V) and Vin=2V..... the answers are: 4.0009857 -0.4990 But i used all the three chances but cudn't figure out the equations, have anybody figured out the equations??

OpenStudy (maroua):

and for Vs=7V, R=8 KOhm k=3.5(mA/√V) and VIN=3.5V, the answers are: 5.24805122253 -0.49990

OpenStudy (maroua):

those are the equations I used: Vout = Vs - ID*R = Vs - R*k*sqrt(Vgs) Vgs = VIN - 2ID*R

OpenStudy (anonymous):

VS=6V, R=10kΩ, and the transistor parameter k=5mAV√. The amplifier is biased with VIN=1V. Can you please help for this?

OpenStudy (anonymous):

Marouna, using your data and formula, i can not get the answers you got, plz explain......

OpenStudy (anonymous):

hello! moukos r u done Q5 question? please tell me.